A support supplies a force 5 kN to a wall as shown below. What is the x component of the moment (in kN-m) produced by this force about an axis in the direction of i+j+k(from the origin)?
1
Expert's answer
2020-07-06T17:25:53-0400

Force is 5 kN


Horizontal Component "i = F_x = 5 cos \theta"


Vertical Component "j =F_y = 5 sin \theta"


 "x" component of the moment (in kN-m)  "= M = F_x \times d = 5 cos \theta \times d (kN-m)"

"d" is the distance of the force form origin


Force about an axis in the direction of "i+j+k = F_x + F_y +F_z = 5cos\theta+5sin\theta+0=5(cos\theta+sin\theta)"